# Diffusion-based Generative Image Outpainting for Recovery of FOV-Truncated CT Images

### v1

@@ Abstract
Field-of-view (FOV) recovery of truncated chest CT scans is crucial for accurate body composition analysis, which involves quantifying skeletal muscle and subcutaneous adipose tissue (SAT) on CT slices. This, in turn, enables disease prognostication. Here, we present a method for recovering truncated CT slices using generative image outpainting. We train a diffusion model and apply it to truncated CT slices generated by simulating a small FOV. Our model reliably recovers the truncated anatomy and outperforms the previous state-of-the-art despite being trained on 87% less data.
@@ 4.1 Quantitative Evaluation
In recovering muscle and SAT areas (Table 1), CT-Palette achieves a significant overall reduction in RMSE compared to the best-performing baseline S-EFOV (56.36% for muscle area, 36.66% for SAT area), despite S-EFOV being pre-trained on a substantially larger dataset.
The box plots in Figure 4 illustrate that the differences in muscle and SAT areas between the untruncated and outpainted slices by CT-Palette follow a Gaussian-like distribution, with an equal tendency for over- or underestimation. CT-Palette’s box plots exhibit a narrow interquartile range and a median close to zero, indicating no significant difference in area between untruncated and outpainted slices. In contrast, S-EFOV consistently underestimates both muscle and SAT areas, likely due to differences between our dataset and the dataset on which S-EFOV was originally trained. Fine-tuning S-EFOV shifts the distribution of SAT area differences toward a greater inclination for overestimation. Lastly, our method achieves the highest Dice score overall, significantly outperforming S-EFOV by 0.31% in muscle and 0.63% in SAT (see supp. material).
@@ Acknowledgments
Appendix 0.A Image Generation Quality (FID)
| Method | Overall (n=997) | T5 (n=339) | T8 (n=348) | T10 (n=310) |
| RFR-Net | 17.505 | 25.277 | 21.571 | 19.471 |
| S-EFOV | 4.809 | 8.106 | 6.416 | 5.405 |
| S-EFOV (FT) | 3.968 | 6.714 | 5.427 | 4.921 |
| CT-Palette (MI) | 3.567 | 5.625 | 5.189 | 5.191 |
Appendix 0.B Muscle and SAT Segmentation Quality (DSC)
| Method | Overall (n=997) | T5 (n=339) | T8 (n=348) | T10 (n=310) | ||||
| Muscle | SAT | Muscle | SAT | Muscle | SAT | Muscle | SAT | |
| Truncated | 0.940 0.083 | 0.865 0.132 | 0.932 0.078 | 0.862 0.114 | 0.933 0.096 | 0.861 0.141 | 0.956 0.070 | 0.874 0.140 |
| RFR-Net | 0.957 0.057 | 0.901 0.078 | 0.963 0.037 | 0.906 0.068 | 0.946 0.072 | 0.894 0.086 | 0.962 0.055 | 0.902 0.078 |
| S-EFOV | 0.976 0.040 | 0.948 0.060 | 0.977 0.029 | 0.944 0.059 | 0.969 0.052 | 0.944 0.066 | 0.982 0.033 | 0.958 0.052 |
| S-EFOV (FT) | 0.974 0.043 | 0.947 0.057 | 0.979 0.024 | 0.947 0.051 | 0.965 0.054 | 0.944 0.059 | 0.978 0.044 | 0.950 0.062 |
| CT-Palette (MI) | 0.979 0.036 | 0.954 0.045 | 0.984 0.019 | 0.957 0.038 | 0.972 0.045 | 0.951 0.050 | 0.980 0.038 | 0.954 0.045 |
Appendix 0.C Qualitative Results of Radiologists’ Study

### v2

@@ Abstract
Field-of-view (FOV) recovery of truncated chest CT scans is crucial for accurate body composition analysis, which involves quantifying skeletal muscle and subcutaneous adipose tissue (SAT) on CT slices. This, in turn, enables disease prognostication. Here, we present a method for recovering truncated CT slices using generative image outpainting. We train a diffusion model and apply it to truncated CT slices generated by simulating a small FOV. Our model reliably recovers the truncated anatomy and outperforms the previous state-of-the-art despite being trained on 87% less data. Our code is available at https://github.com/michelleespranita/ct_palette.
@@ 4.1 Quantitative Evaluation
In recovering muscle and SAT areas (Table 1), CT-Palette achieves a significant overall reduction in RMSE compared to the best-performing baseline S-EFOV (56.36% for muscle, 36.66% for SAT when using multiple inference; 44.70% for muscle, 20.38% for SAT when using single inference), despite S-EFOV being pre-trained on a substantially larger dataset.
The box plots in Figure 4 illustrate that the differences in muscle and SAT areas between the untruncated and outpainted slices by CT-Palette follow a Gaussian-like distribution, with an equal tendency for over- or underestimation. CT-Palette’s box plots exhibit a narrow interquartile range and a median close to zero, indicating no significant difference in area between untruncated and outpainted slices. In contrast, S-EFOV consistently underestimates both muscle and SAT areas, likely due to differences between our dataset and the dataset on which S-EFOV was originally trained. Fine-tuning S-EFOV shifts the distribution of SAT area differences toward a greater inclination for overestimation. The Wilcoxon signed-rank test shows no significant difference in the muscle/SAT area distribution between ground-truth images and those outpainted by our CT-Palette, while other models show a significant difference (see supp. material).
Lastly, our method achieves the highest Dice score overall (0.979 for muscle, 0.954 for SAT when using multiple inference), significantly outperforming S-EFOV by 0.31% in muscle and 0.63% in SAT (see supp. material).
@@ 5.0.1 Acknowledgements
[…]
The authors have no competing interests to declare that are relevant to the content of this article.